### v1

# To Trust or Not to Trust: On Calibration in ML-based Resource Allocation for Wireless Networks

Abstract
In the next generation communications and networks, machine learning (ML) models are expected to deliver not only highly accurate predictions, but also well-calibrated confidence scores that reflect the true likelihood of correct decisions. In this paper, we study the calibration performance of an ML-based outage predictor within a single-user, multi-resource allocation framework. We begin by establishing key theoretical properties of this system’s outage probability (OP) under perfect calibration. Importantly, we show that as the number of resources grows, the OP of a perfectly calibrated predictor approaches the expected output conditioned on it being below the classification threshold. In contrast, when only a single resource is available, the system’s OP equals the model’s overall expected output. We then derive the OP conditions for a perfectly calibrated predictor. These findings guide the choice of the classification threshold to achieve a desired OP, helping system designers meet specific reliability requirements. We further demonstrate that post-processing calibration cannot improve the system’s minimum achievable OP, as it does not introduce additional information about future channel states. Additionally, we show that well-calibrated models are part of a broader class of predictors that necessarily improve OP. In particular, we establish a monotonicity condition that the accuracy-confidence function must satisfy for such improvement to occur. To demonstrate these theoretical properties, we conduct a rigorous simulation-based analysis using post-processing calibration techniques, namely, Platt scaling and isotonic regression. As part of this framework, the predictor is trained using an outage loss function specifically designed for this system. Furthermore, this analysis is performed on Rayleigh fading channels with temporal correlation captured by Clarke’s 2D model, which accounts for receiver mobility. Notably, the outage investigated refers to the required resource failing to achieve the transmission capacity requested by the user.
Index Terms:
Custom loss function, Isotonic regression, Machine learning, ML calibration, Outage prediction, Outage loss function, Outage probability, Post-processing calibration, Platt scaling, Resource allocation, Reproducible AI, Trustworthy AI.
I Introduction
How should we interpret a neural network’s (NN) confidence score? If a NN assigns a confidence of 0.4 to the question “Will this communication channel fail in the future?” what does that really mean? What if it instead returns 0.000001? These questions highlight the importance of calibration—the extent to which a network’s confidence estimates align with the true probability of an event occurring.
A well-calibrated NN allows us to treat its outputs as reliable probabilities, making its predictions more actionable in real-world decision-making. For instance, suppose a user can tolerate a channel failure rate of at most 1 in 100 transmissions. They could consult the NN before using a channel, and if the model predicts a failure probability below 0.01, the channel is deemed acceptable; otherwise, they should seek an alternative. However, this approach is only effective if the NN is well-calibrated, particularly around the critical threshold of 0.01.
Accurate probability estimates are essential for informed decisions, yet NNs often produce confidence scores that do not accurately reflect true probabilities, a phenomenon known as poor calibration. When a model is miscalibrated, it may systematically overestimate or underestimate the likelihood of an event, leading to unreliable decision-making. Proper calibration aligns predicted probabilities with actual outcomes, enhancing both the reliability and usability of machine learning (ML)-driven predictions [1].
One of the most common ways to evaluate calibration is through reliability diagrams, which compare predicted confidence levels with observed accuracy (see Fig. 1(top)). Interestingly, in communication systems, the failure probabilities that users care about can span several orders of magnitude, from rare events to more frequent occurrences, making standard calibration visualizations on linear scales inadequate. To address this, logarithmic binning[2] provides a more effective way to assess calibration across broad probability ranges, offering a clearer picture of how well NNs perform at low probability events.
Improving calibration is particularly critical in ML-driven wireless networks, where miscalibrated models can result in inefficient resource allocation, increased failure rates, or unnecessary precautionary measures. This issue becomes even more significant in ultra-reliable low-latency communications (URLLC) applications, such as industry automation, where accurately predicting rare outage events is essential for maintaining safety and system efficiency. Motivated by these concerns, we propose applying post-processing calibration methods to improve the reliability of an ML-based outage predictor. We examine how its calibration properties influence resource allocation errors in a single-user, multi-resource allocation framework. In particular, we focus on the classification threshold of this predictor, which determines how calibrated confidence scores are used to predict outages and helps achieve a target system outage probability (OP), which serves as our metric for resource allocation error.
I-A Related Work
I-A1 ML for Wireless Systems
ML has become an essential tool for ensuring reliable connectivity and efficient resource allocation in dynamic wireless networks. In particular, deep learning (DL) has been widely applied to tasks such as blockage prediction and link quality assessment [3, 4, 5, 6, 7]. For example, in [3], a deep NN was trained to map user positions and traffic demands to blockage status and optimal beam index, achieving a blockage prediction accuracy of 90% in millimeter wave (mmWave) systems. The authors of [4] developed a meta-learning based framework for mmWave and terahertz systems, where a recurrent NN was trained using a few data samples to predict potential blockages. In [5], models were trained on mmWave received power sequences to identify pre-blockage signatures. In [6], beam sequence data was processed using a gated recurrent unit network to forecast line-of-sight blockages. Additionally, [7] proposed a NN that learned to predict both the optimal mmWave beam and the blockage status using sub-6 GHz channel information.
Beyond blockage prediction, DL has also been employed for power allocation in massive multiple-input multiple-output (MIMO) systems [8] and channel estimation [9, 10, 11, 12]. For instance, [8] demonstrated a significant reduction in the complexity and processing time of the optimization process for massive MIMO systems; [9] used a convolutional NN for wideband channel estimation in mmWave massive MIMO systems; [10] and [11] used long short-term memory (LSTM) networks to make channel predictions in body area networks and future signal strength, respectively; and [12] explored DL techniques to predict future received signal strength variations in device-to-device (D2D) communications channels. DL has also found extensive use in traffic prediction. For example, [13] proposed a hybrid DL model, while [14] applied a recurrent NN for spatiotemporal traffic prediction in cellular networks; [15] compared the performance of statistic learning and DL in user traffic prediction; and [16] developed a DL architecture and applied transfer learning among different scenarios for traffic prediction.
Further expanding its utility, numerous studies have used ML techniques, such as deep reinforcement learning (DRL), for a wide range of wireless communication tasks, including mobile networking [17, 18, 19], spectrum sharing in cognitive radio networks [20], scheduler design [21, 22, 23], resource allocation [24, 25, 26, 27, 28, 29, 30, 31, 32], and energy efficiency optimization in D2D enabled heterogeneous networks [33]. For example,[17] addressed internet congestion by dynamically adjusting sender data rates, while[21] applied DRL to select scheduling policies that minimize packet delays and drop rates. In [24], the authors proposed an optimal classifier for an ML-assisted resource allocation system, [25] minimized long-term system power consumption in Fog radio access network, and [26] proposed a decentralized DRL-based framework for resource allocation in vehicle-to-vehicle communications, effectively supporting both unicast and broadcast modes. Similarly, [27] minimized average latency in mobile edge computing (MEC) Internet-of-Things networks. In a related context, approaches such as Quantum DRL have also been used to support joint resource allocation and task offloading in MEC [34] and to enhance computational learning speed [35]. Additionally, several studies have proposed ML-driven strategies for resource allocation in URLLC systems[36, 37], as well as for managing the coexistence of URLLC and enhanced mobile broadband services in 5G New Radio networks [38, 39, 40].
I-A2 Custom Loss Functions for ML-Driven Wireless Systems
Most ML-based communication models highlighted previously, rely on traditional loss functions, such as binary cross-entropy (BCE) and mean squared error (MSE), to optimize model performance. While these loss functions are effective for general classification and regression tasks, they do not always align with the specific requirements of wireless networks [41]. In particular, conventional training objectives often fail to capture communication-specific constraints, such as OPs, resource efficiency, and latency guarantees. This misalignment can lead to suboptimal performance, prompting research into more specialized optimization approaches.
To address these limitations, custom loss functions have been designed to incorporate domain knowledge and better reflect communication system objectives [42]. Such loss functions have been successfully applied in hybrid beamforming for MIMO systems[43], optimizing reflection coefficients in reconfigurable intelligent surfaces[44], and resource allocation in cell-free massive MIMO networks[45] and URLLC[37]. In addition,[46] introduced the outage loss function (OLF) to specifically minimize OPs during model training. Other applications include wireless channel estimation, where they preserve low-rank properties of the estimated channel matrix [47], cellular traffic prediction [48], received signal strength indicator (RSSI)-based wireless indoor localization [49], and statistical federated learning for beyond 5G network slicing [50].
I-A3 ML Calibration
Regardless of the loss function used, calibration remains a key challenge in ML-driven communication systems. To address this issue, Bayesian learning [51, 52] models epistemic uncertainty through parameter distributions to improve probabilistic predictions. It has been applied to a range of communication tasks, including network resource allocation [53, 54, 55], massive MIMO detection [56, 57, 58], and channel estimation [59]. While exact Bayesian learning offers formal calibration guarantees, these rely on strong assumptions. Specifically, they assume that the model is well specified, meaning the NN has enough capacity to represent the ground-truth data generation process, and that its predictive distribution for continuous outputs is unimodal. These assumptions may not be satisfied in cases where sufficient data is not available. In [60], the authors explored the application of robust Bayesian learning to wireless communication systems, focusing on applications such as automated modulation classification, RSSI-based localization, as well as channel modeling and simulation. However, this approach does not offer formal calibration guarantees. Conformal prediction [61] is a general framework for constructing set predictors that offer formal, distribution-free calibration guarantees and has been applied to problems such as symbol demodulation and modulation classification [62]. More recently, it has been explored as a tool to support uncertainty quantification through pre-deployment calibration, online monitoring, and post-deployment diagnostic analysis [63]. Another widely used approach is post-processing calibration, which adjusts confidence scores using a validation set. Common techniques include Platt scaling, isotonic regression, beta calibration for binary classification, and temperature or Dirichlet scaling for multiclass settings [64, 65, 66, 67, 68, 69, 70].
I-B Contributions
Our contribution focuses on studying the calibration of the outage predictor proposed in [46] for an ML-assisted resource allocation system, providing a more precise understanding of its role in resource allocation. Specifically, we summarize our key contributions as follows:
  1. 1.
    We demonstrate that in the infinite resource limit, the OP of the system, when the perfectly calibrated model is used equals, its expected confidence, conditioned on it being less than its classification threshold.
  2. 2.
    We show that when only one resource is available, and the model is perfectly calibrated, the model’s expected output is equal to the resource’s underlying OP.
  3. 3.
    Our theoretical analysis shows that, for well-calibrated models, there is a straightforward approach for choosing a classification threshold to achieve a desired system OP. For example, when a model is well calibrated and the number of available resources is sufficiently large, system designers need only choose a classification threshold equal to their desired OP. The resulting system OP will then be less than their selected threshold.
  4. 4.
    We show that well-calibrated models form a special case of a more general class of model that necessarily improves OP. Specifically, we establish a monotonicity condition that accuracy-confidence functions must satisfy in order to improve system OP.
  5. 5.
    We show that post-processing calibration cannot improve the system’s minimum achievable OP. This follows from the data processing inequality, which ensures that such calibration methods do not introduce additional information about future channel states.
  6. 6.
    A rigorous simulation-based analysis in a Rayleigh fading environment using post-processing calibration techniques is conducted, demonstrating the theoretical findings above.
  7. 7.
    Finally, we present a preliminary proof of concept demonstrating the application of the resource allocation technique and the outage loss function used in this work and in [71]. This demonstration is intended for a broader audience.
This work builds on [2, 72] and presents a significantly more comprehensive study of the outage predictor calibration within this system. We analyze how calibration affects the classification threshold, enabling more effective resource selection, while also incorporating the additional contributions outlined above. For reproducibility purposes, we have made our code publicly available [^1] https://github.com/ML4Comms/greedy-resource-allocation-outage-classification [71]..
The remainder of this paper is organized as follows: Section II revisits the resource allocation strategy incorporating an outage predictor, as proposed in [46]. It then introduces the OLF used for its training, before providing an overview of calibration concepts. Section III discusses the key theoretical results of this work. Section IV begins with a description of data generation process, outlines the essential metrics for evaluating calibration performance, and provides an overview of the post-processing calibration techniques implemented in this paper. Section V presents the numerical results that demonstrate the findings established in Section III, and discusses the proof of concept for the resource allocation technique introduced in Section II. Finally, Section VI concludes the paper with some finishing remarks.
II System Model
In this section, we revisit the single-user, multi-resource system model from [46] and restate the essential formulations to provide the necessary context for this study. We focus on the ML-based outage predictor utilized by the user and examine how its calibration properties influence the resource allocation framework.
II-A Channel Model for a Resource
Each resource $j\in\mathcal{R}$ has a fluctuating channel state $h_{j}(t)\in~{}\mathbb{C}$, with channel states assumed to be independent and identically distributed (i.i.d.) across distinct resources. Furthermore, the correlation between $h_{j}(t)$ and $h_{j}(t+l)$ decreases as $l\to\infty$. The user is equipped with an ML-based outage predictor that assists in selecting the appropriate resource by learning these temporal correlation patterns. For $k,l\in\mathbb{N}$, the vectors of past and future channel states at time $t$ are given by $\accentset{\leftarrow}{H}_{j}^{k}{(t)}\triangleq\left[h_{j}{(t-k+1)},\cdots,h_{j}{(t)}\right]^{T}$, and $\accentset{\rightarrow}{H}_{j}^{l}(t)\triangleq\left[h_{j}(t+1),\cdots,h_{j}(t+l)\right]^{T}$, respectively.
The capacity $C\left({\accentset{\rightarrow}{H}_{j}^{l}}{(t)}\right)\in\mathbb{R}^{+}$ measures resource $j$’s ability to support communication from time $t+\!1$ to $t+\!l$. For a quasi-static Gaussian channel it is given by [73, eq. (5.80)]:
$\!C\!\left(\!{\accentset{\rightarrow}{H}_{j}^{l}}{(t)}\!\right)\!=\!\sum_{i=1}^{l}\log_{2}\left(1+\mathtt{SNR}\left|h_{j}\left(t+i\right)\right|^{2}\right)~{}~{}~{}~{}\rm{\!bits\!/s/\!Hz},$ (1)
where $\mathtt{SNR}$ is the average signal-to-noise ratio per sample. If the user’s requirements, defined by a key performance indicator (KPI) such as transmission rate, exceed this capacity, the resource is considered to be in outage; otherwise, it is deemed satisfactory. The required KPI level is specified by a rate threshold, $\gamma_{th}$.
The OP for a single resource $j$ can be expressed as
$P_{j}(\gamma_{th})=\mathbb{P}\left[C\left({\accentset{\rightarrow}{H}_{j}^{l}}{(t)}\right)<\gamma_{th}\right],$ (2)
where $\mathbb{P}[\cdot]$, represents the ground-truth distribution of the observation state sequences $\accentset{\rightarrow}{H}_{j}^{l}{(t)}$.
II-B ML-based Resource Allocation
This work focuses on a calibrated ML-based predictor, unlike [46], which considers the general case where the model may or may not be well-calibrated. The predictor’s output $Q\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)};\Theta\right)\in[0,1]$, determined by the parameters $\Theta$, provides an estimate of the conditional probability:
$\mathbb{P}\left[C\left(\accentset{\rightarrow}{H}_{j}^{l}{(t)}\right)<\gamma_{th}\mid\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right],$
where the conditioning is over the information $\accentset{\leftarrow}{H}_{j}^{k}{(t)}$ available at the time of prediction. An outage is predicted if $Q\!\left(\!\accentset{\leftarrow}{H}_{j}^{k}{(t)};\Theta\right)\!>\!q_{th}$, where $q_{th}$ is the predictor’s classification threshold.
The user relies on an ML-based outage predictor for resource allocation, where resources are indexed from $j=1$ to $|\mathcal{R}|$. Each time the predictor predicts an outage event, the index is incremented by 1. The predictor uses the input vector $\accentset{\leftarrow}{H}_{j}^{k}{(t)}$ for each resource $j$, producing an output in $[0,1]$ that classifies whether the upcoming $l$ channel samples $\accentset{\rightarrow}{H}_{j}^{l}{(t)}$ will support outage-free communication. If no resource meets this condition, the model selects the final resource.
In order to quantify the calibration quality of the predictor, we adopt the standard characterization in terms of reliability diagrams. A reliability diagram is a visual representation of model calibration, plotting the observed accuracy as a function of confidence [1]. In this context, the accuracy-confidence function $A_{j}(q)$ describes the probability of an outage event given the predictor provides an outage confidence of $q$. For our outage predictor $Q\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)$, the accuracy-confidence function is formally defined as
$A_{j}(q)\triangleq\mathbb{P}\left[C\left(\accentset{\rightarrow}{H}_{j}^{l}{(t)}\right)<\gamma_{th}\,\middle|\,Q\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)};\Theta\right)=q\right].$ (3)
Figure 1: Reliability diagram showing the accuracy-confidence function $A_{j}(q)$ (top), and the confidence distribution $F_{j}(q)$ (bottom).
The accuracy-confidence function gauges the degree to which a predictor’s confidence level matches the true predictive accuracy for any fixed confidence level $q$. This does not provide information on how frequently the predictor returns a confidence level $q$. That information is given by the confidence distribution $F_{j}(q)$, defined as the cumulative distribution function
$F_{j}(q)=\mathbb{P}\left[Q\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)\leq q\right],$ (4)
which gives the fraction of past-sample histories $\accentset{\leftarrow}{H}_{j}^{k}{(t)}$ for which the predictor returns a confidence level less than or equal to $q$.
Illustrative plots of the functions $A_{j}(q)$ and $F_{j}(q)$ are shown in Fig. 1. The top figure shows three accuracy-confidence curves corresponding to the extreme cases: perfect calibration $\left(A_{j}(q)=q\right)$, strong under-confidence $\left(A_{j}(q)>q\right)$, and strong over-confidence $\left(A_{j}(q)<q\right)$, for all $q\in[0,1]$. The bottom figure shows two examples of confidence distributions, where lower $F_{j}(q)$ indicates a model that tends to produce larger fraction of confident decisions.
Since the predictor is generally incorrectly calibrated, we have the non-equality
$A_{j}(q)\neq q$ (5)
The goal of this paper is to study the effect of applying calibration strategies to improve the calibration properties of predictive models, so that
$A_{j}(q)\approx q.$ (6)
Let $Q^{\star}\in\{Q,Q^{c}\}$, where $Q^{c}$ denotes the calibrated predictor. Similar to [1], we assume that the calibrated predictor be a post-processing of the original scores:
$Q^{c}\;=\;\mathcal{C}\bigl{(}Q;\mathcal{D}\bigr{)}\;=\;r\circ Q,$ (7)
where the recalibration map
$r:[0,1]\;\longrightarrow\;[0,1]$ (8)
is learned solely from a held-out calibration dataset $\mathcal{D}$. Here, $\mathcal{C}\bigl{(}Q;\mathcal{D}\bigr{)}$ represents the calibration procedure, and $r\circ Q$ represents function composition i.e., the calibrated predictor is obtained by applying $r$ to the output of $Q$. Since $Q^{c}\!\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)$ is obtained by (possibly stochastic) processing of $Q\!\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)$, the data processing inequality implies
$\!I\!\left(\!C\!\left(\!\accentset{\rightarrow}{H}_{j}^{l}{(t)}\!\right);\!Q^{c}\!\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\!\right)\!\right)\;\!\!\leq\;I\!\left(C\left(\accentset{\rightarrow}{H}_{j}^{l}{(t)}\!\right);\!Q\!\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\!\right)\!\right)\!,$ (9)
where $I(\cdot,\cdot)$ is the mutual information [74]. This ensures that no additional information about $C\left(\accentset{\rightarrow}{H}_{j}^{l}{(t)}\right)$ is injected during calibration. Next, we describe the resource allocation strategy used in this study.
II-C Greedy Resource Allocation
We adopt the resource allocation strategy from [46], in which the user scans the available resources $\mathcal{R}$ and selects a resource $j^{\prime}$ based on whether the outage predictor $Q^{\star}\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)$ indicates successful communication. The predictor sequentially scans resources =$\mathcal{R}=\{1,2,\dots,|\mathcal{R}|\}$, stopping once it predicts successful communication for a given resource. If no resource meets this condition, the final resource is selected.
Formally, let
~$\tilde{\mathcal{R}}({q_{th}})=\left\{j\in\mathcal{R}~{}\text{s.t.}~{}Q^{\star}\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right);\Theta)\leq q_{th}\right\},$ (10)
represent the subset of resources where the outage predictor predicts no outages. Then the greedy allocation scheme selects
$j^{\prime}(q_{th})=\begin{cases}\min_{j\in\tilde{\mathcal{R}}(q_{th})}~{}j~{}~{}&\text{if }\tilde{\mathcal{R}}({q_{th}})\neq\emptyset,\\ |\mathcal{R}|&\text{otherwise}.\end{cases}$ (11)
Let $P^{\star}_{|\mathcal{R}|}(\gamma_{th},q_{th})\in\{P_{|\mathcal{R}|}(\gamma_{th},q_{th}),\;P^{c}_{|\mathcal{R}|}(\gamma_{th},q_{th})\}$ denote the system’s OP when using either the uncalibrated predictor $Q$ or the calibrated predictor $Q^{c}$, respectively. These are formally defined as:
$P^{\star}_{|\mathcal{R}|}(\gamma_{th},q_{th})\triangleq\mathbb{P}\left[C\left(\accentset{\rightarrow}{H}_{j^{\prime}(q_{th})}^{l}(t)\right)<\gamma_{th}\right],$ (12)
where $j^{\prime}(q_{th})$ is the selected resource index according to (11). Similarly, in the infinite-resource limit, as derived in [46], the OP is given by:
$\!\!\!P^{\star}_{\infty}(\gamma_{th},q_{th})\!=\mathbb{P}\!\left[C\left(\accentset{\rightarrow}{H}_{j}^{l}(t)\!\right)\!<\gamma_{th}\middle|Q^{\star}\!\left(\accentset{\leftarrow}{H}_{j}^{k}(t)\!\right)\leq q_{th}\right]\!,$ (13)
with $Q^{\star}\in\{Q,Q^{c}\}$ as before.
Following [46, eq. (13)] and using the law of total probability, the OP in (12), expressed in terms of the predictor’s confidence distribution function and accuracy-confidence function is given by:
$P^{\star}_{|\mathcal{R}|}(\gamma_{th},q_{th})$ $=P_{j}(\gamma_{th})\left(1-F_{j}(q_{th})\right)^{|\mathcal{R}|-1}$
$+\!\mathbb{E}\!\left[A_{j}\left(Q^{\star}\!\left(\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)\right)\mid\!Q^{\star}\left(\!\accentset{\leftarrow}{H}_{j}^{k}{(t)}\right)\!\leq q_{th}\right]$
×$\times\left(1-(1-F_{j}(q_{th}))^{|\mathcal{R}|-1}\right).$ (14)
Furthermore, we say that a predictor is useful at a threshold level $q_{th}$ if the inequality
$P^{\star}_{|\mathcal{R}|}(\gamma_{th},q_{th})<\underset{j\in\mathcal{{R}}}{\text{min}}~{}P_{j}$ (15)
holds. This indicates that the predictor allows the selection of a resource (11) that supports a probability of outage lower than the minimum OP $\min_{j\in\mathcal{R}}~{}P_{j}$ achievable through random selection of a resource.
The OLF, customized for this resource allocation system, is defined in [46, eq. (29)]. This loss function is utilized to train the outage predictor, enabling the generation of the results presented in Section V. For brevity, we drop the dependence of the sequences $\accentset{\leftarrow}{H}_{j}^{k}{(t)}$ and $\accentset{\rightarrow}{H}_{j}^{l}{(t)}$ on $t$, as well as the dependence on $j$ for $P\triangleq P_{j}$, accuracy-confidence function $A(q)\triangleq A_{j}(q)$, and confidence distribution $F(q)\triangleq F_{j}(q)$. Additionally, where appropriate, we drop the dependence of the predictor’s confidence $Q^{\star}(\cdot)$ and of $C(\cdot)$ on $\accentset{\leftarrow}{H}_{j}^{k}{(t)}$ and $\accentset{\rightarrow}{H}_{j}^{l}{(t)}$, respectively. We denote by $Q^{\star}$ a random variable with the same distribution as any confidence level $Q^{\star}$ assigned by the predictor to a resource $j$.
Before moving on to the next section, we reiterate our problem statement here. This paper builds on [2], which analyzed the general case where the outage predictor may or may not be well calibrated for an ML-assisted resource allocation system. In such cases, the predictor’s confidence scores may not accurately reflect the true probability of outages. Central to our study is the characterization of the system’s OP under perfect calibration which helps guide the choice of the classification threshold required to achieve a target OP when sufficient resources are available. We then apply post-processing calibration techniques to validate our theoretical findings.
III On the Impact of Calibration
In this section, we present the main results of this work, detailing how the output of the calibrated predictor relates to the system’s outage performance metrics.
III-A Main Results
Our analysis reveals that perfectly calibrated predictors satisfy key properties related to the system’s OP. Specifically, for $|\mathcal{R}|\to\infty$, the OP equals the expected confidence conditioned on $Q^{c}\leq q_{th}$; in the single resource case, it equals the overall expected confidence. These results help identify an appropriate $q_{th}$ to meet target performance results, providing guidance for designing wireless systems with specific reliability levels. This is captured in the following theorem.
Theorem 1.
Consider the resource allocation system described in Section II. Then, for a perfectly calibrated predictor, we have:
$P$ $=\mathbb{E}[Q^{c}],$ (16)
$\text{and}\qquad P^{c}_{\infty}(\gamma_{th},q_{th})$ $=\mathbb{E}[Q^{c}\mid Q^{c}\leq q_{th}],$ (17)
so that we can express:
$\noindent\!\!P^{c}_{|\mathcal{R}|}(\gamma_{th},q_{th})\!$ $=\mathbb{E}[Q^{c}](1-\!F(q_{th}))^{|\mathcal{R}|-1}$
$\hskip 1.00006pt+\mathbb{E}[Q^{c}\!\mid Q^{c}\!\leq q_{th}]\!\left(\!1-\!(1-\!F(q_{th}))^{|\mathcal{R}|-1}\!\right)\!.$ (18)
Proof:
See Appendix A-A
Theorem 1 reveals that for a perfectly calibrated predictor, the OP is determined by the predictor’s expected confidence, either conditioned on being below $q_{th}$ (in the infinite-resource case) or across all predictions (in the single-resource case). Moreover, (17) tells us that when the number of resources is sufficiently large, one should select a value of $q_{th}$ such that the conditional expected OP of the model is equal to their desired OP. Using Theorem 1, we can further establish simpler results on the relationship between the system’s OP and the classification threshold $q_{th}$ in the well-calibrated scenario.
Corollary 1.
For a perfectly calibrated predictor, the OP satisfies:
$P^{c}_{\infty}(\gamma_{th},q_{th})\leq q_{th},$ (19)
$P^{c}_{|\mathcal{R}|}(\gamma_{th},q_{th})$ $\leq\mathbb{E}[Q^{c}](1-F(q_{th}))^{|\mathcal{R}|-1}$
$+q_{th}\left(1-(1-F(q_{th}))^{|\mathcal{R}|-1}\right).$ (20)
This corollary further simplifies the process of selecting an appropriate $q_{th}$. For example, (19) tells us that if the model is well-calibrated, all one needs to do is select a $q_{th}$ equal to their desired OP, and provided there are sufficient resources, they should achieve a system OP that is better than their desired OP.
Note that although calibration improves the alignment between predicted confidence and actual accuracy, it does not increase the informativeness of the predictor, as it does not introduce new information. As a result, post-processing calibration cannot improve the system’s minimum achievable OP, as formalized in the following theorem.
Theorem 2.
For any post-processing calibration procedure satisfying (7), applying this calibration will not improve the system’s OP performance. i.e.,
$\min_{q_{th}}P^{c}_{|\mathcal{R}|}(\gamma_{th},q_{th})\geq\min_{q_{th}}P_{|\mathcal{R}|}(\gamma_{th},q_{th})$ (21)
Proof:
See Appendix A-B
Figure 2: Illustration of Theorem 3, where (a) a predictor is useful at $q_{th}=0.5$, satisfying (22), and (b) a predictor is not useful at the same $q_{th}$.
We now come to the final theoretical result of our work which introduces the concept of monotonicity that the accuracy-confidence function must satisfy for the predictor to be useful (i.e., satisfying (15)).
Theorem 3.
Given a threshold level $q_{th}$, a predictor is useful (i.e., it satisfies (15)) if the inequality $\mathbb{E}\left[A(Q^{\star})|Q^{\star}\leq q_{th}\right]$ $<$ $P$ is satisfied. Furthermore, $\mathbb{E}\left[A(Q^{\star})|Q^{\star}\leq q_{th}\right]<P$ is satisfied if and only if
$\mathbb{E}\left[A(Q^{\star})|Q^{\star}\leq q_{th}\right]<\mathbb{E}\left[A(Q^{\star})|Q^{\star}>q_{th}\right].$ (22)
Moreover, at the critical point defined by:
$\mathbb{E}[A(Q^{\star})\mid Q^{\star}\leq q_{th}]=\mathbb{E}[A(Q^{\star})\mid Q^{\star}>q_{th}]$ (23)
there is no net change in the system’s outage performance.
Proof:
See Appendix A-C
This theorem shows us that the conditional expectation of the predictor’s accuracy-confidence function should be smaller when conditioned on values less than or equal to $q_{th}$ than when conditioned on values greater than $q_{th}$. A trivial class of accuracy-confidence functions that achieve this are those that are monotonically increasing. As an example, if $A(q)=q$ for all values of $q$, as is the case for perfectly calibrated predictors, then $A$ is a monotonically increasing function and so the corresponding predictor necessarily improves the OP. There may be examples of strongly over-confident and strongly under-confident predictors that also satisfy this monotonicity property (see Fig.1 (top)). Importantly, $A(q)$ being monotonically increasing is a sufficient but not necessary condition required for a predictor to be useful. For example, it is possible for $A(q)$ to be non-monotonic in the regions $[0,q_{th})$ and $[q_{th},1]$ whilst also satisfying $A(q)>A(q^{\prime})$ for $q^{\prime}\in[0,q_{th})$ and $q\in[q_{th},1]$. Such a function would also satisfy (22) (see Fig.2 (a)). Conversely, we can also see that if a predictor’s accuracy-confidence function is monotonically decreasing, then it will necessarily hinder the system’s OP (see Fig.2 (b)). Again, using a similar example as above, this is a sufficient but not necessary property of a predictor that hinders performance.
Furthermore, this theorem tells us that at a critical point, defined by the condition that the average accuracy for instances where $Q^{\star}\leq q_{th}$ equals that for $Q^{\star}>q_{th}$, the system’s outage performance exhibits no overall change.
IV Methodology
This section describes the data-generation process and outlines the key metrics used to evaluate calibration performance of the outage predictor. It also provides an overview of the post-processing calibration techniques implemented in this study, along with relevant implementation details.
IV-A Data Generation
The training/testing dataset for the channel is synthetically obtained as follows. We start by generating a zero-mean complex Gaussian vector given by: $G[0]=[g_{1},g_{2},\dots,g_{v}]$ at $t=0$, with $v=1024$, where each element has a variance of $\frac{1}{v}$. To simulate the motion of a mobile user, at each time step $t$, we apply a per-element, independent and uniformly distributed phase shifts $\theta_{it}\sim\text{Unif}\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$. At time $t=1$, we have:
$G[1]=\Bigg{[}e^{i\xi\sin\theta_{11}}g_{1},e^{i\xi\sin\theta_{21}}g_{2},\ldots,e^{i\xi\sin\theta_{v1}}g_{v}\Bigg{]},$ (24)
and so on. In the frequency domain, we have:
$\!\mathcal{H}[1]\!=\!\frac{1}{\!\sqrt{v}}\!\Bigg{[}$ $\!\sum_{m=0}^{v-1}\!e^{i\xi\!\sin\theta_{\!m1}}\!g_{m}\!e^{\!-i2\!\pi\!m/v},\!\ldots,\sum_{\!m=0}^{\!v-1}\!e^{i\xi\!\sin\theta_{\!m1}}\!g_{m}\!e^{\!-i2\!\pi\!m}\!\Bigg{]}.$ (25)
Let $h_{i}(t)$ denote the $i$-th element of $\mathcal{H}[t]$. The channel for resource $j$ at time $t$ is modeled as:
$y_{j}[t]=h_{i}[t]x_{j}[t]+w_{j}[t],$ (26)
where $x_{j}[t]$ represents a unit variance signal term, $w_{j}[t]$ represents a unit variance zero mean complex Gaussian noise term, and $h_{i}[t]$ are stationary and zero mean complex Gaussian variables. Furthermore, the autocorrelation between successive channel samples is given by:
$\mathbb{E}[h_{j}(t)h_{j}(t+1)]$ $=\mathbb{E}\left[\frac{1}{v}\sum_{m=0}^{v-1}e^{i\xi\sin\theta_{m1}}\lVert g_{m}\rVert^{2}\right]$
$=\mathbb{E}\lVert g_{m}\rVert^{2}\mathrm{J_{0}\left(\xi\right)},$ (27)
where $\mathrm{J_{0}(\cdot)}$ is the zeroth-order Bessel function of the first kind, and (27) follows from the characteristic function of the uniform distribution. This model is in-line with the widely accepted Clarke’s 2D model [^2] Additionally, channel data generated using Clarke’s 3D model [75] was used, and the results discussed in the subsequent subsections apply equally to this dataset. [76].
TABLE I: DQN-LSTM Model Architecture and Hyperparameters
Architecture Specification
Layer 1: LSTM layer 32 hidden units
Layer 2: Dense layer 10 units with PReLU activation
Layer 3: Dense layer 1 unit with sigmoid activation
Hyperparameter Values
Hidden units 32
Epochs 30
Epoch size 150
Input sequence length ($k$) 100
Output sequence length ($l$) 10
Output dimension 1
Learning rate 0.001
Discount factor 0.9
Epsilon Decaying strategy
Rewards
1 (correct prediction);
-1 (otherwise)
IV-B Training and Evaluation
We utilize a Deep Q-Network (DQN) incorporating an LSTM layer, following the implementation in [77], and train it on the channel data using both the OLF from [46] ($\alpha=10$) and the BCE loss function. The predictor is developed using TensorFlow’s Keras API and trained with the ADAM optimizer. Table I details the architecture and hyperparameters of the DQN-LSTM predictor employed in our experiments. We train our predictor on a dataset comprising 4,500 × (number of resources) samples. For a 4-resource system, this corresponds to 18,000 training samples, with the dataset size scaling proportionally (e.g., 27,000 samples for a 6-resource system). The training and validation datasets are identical in size, and testing is performed on 11,000 instances of the $|\mathcal{R}|$-resource system to achieve statistically robust results. The predictor is retrained 10 times, and the average performance is used as a single data point for analysis. Each experiment generates sequences of $k+l$ channel samples per resource, where $k=100$, $l=10$, and $\xi=0.1$ radians. The first $k$ samples are provided as input to the DQN-LSTM model, while the last $l$ samples, $\accentset{\rightarrow}{H}_{j}^{l}{(t)}$, are used to determine whether communication can be supported at a specified rate. On a desktop computer equipped with a 12th Gen Intel(R) Core(TM) i9-12900HK, 2500 MHz, 14 Core(s) and 20 Logical processor(s), the training of a single DQN-LSTM model takes approximately 5 minutes.
The reliability diagrams presented in Section V are constructed using logarithmic binning as defined in [2, Eq. (7)]:
$\left(0,1\right]=\bigcup_{n=0}^{\infty}\left[b^{-(n+1)},b^{-n}\right],$ (28)
where $b=\sqrt{10}$. These reliability plots represent the average calibration performance of our outage predictor over 10 runs, with each run evaluating calibration on 10,000 samples.
To empirically demonstrate the key properties of the OP established under perfect calibration in Section III, we apply two widely used post-processing calibration techniques: Platt scaling and isotonic regression. Their effectiveness is evaluated using standard calibration metrics. Specifically, we use Negative log-likelihood (NLL) [1] and Brier score (BS) [78, 79], which assess how well predicted probabilities align with actual outcomes for individual predictions. NLL, also known as cross-entropy loss [80], measures the alignment of predicted probabilities with true labels [66]. The BS captures the mean squared difference between predicted probabilities and true outcomes, with values $\in[0,1]$. For both the metrics, lower values indicate better calibration.
We now briefly describe the calibration techniques implemented using widely available Python libraries, including scikit-learn for logistic and isotonic regression and SciPy for mathematical computations, as detailed below.
  1. 1.
    Platt Scaling: This is a popular method for calibrating binary classifiers due to its simplicity and effectiveness. It fits a logistic regression model to the output logits (raw outputs). For each sample $i$, the calibrated probability is given by:
    $\hat{q}_{i}=\sigma(mz_{i}+n),$ (29)
    where $z_{i}$ is the logit, $m$ and $n$ are parameters learned during calibration, and $\sigma(\cdot)$ is the sigmoid function. Since the DQN-LSTM outputs probabilities through a sigmoid activation, we first convert them to logits using the logit function:
    $\texttt{logit}(p)=\ln\left(\frac{p}{1-p}\right),$ (30)
    where $p$ is the predicted probability. These logits serve as inputs to scikit-learn’s LogisticRegression, which fits the parameters $m$ and $n$ by minimizing the NLL on a validation set using the lbfgs solver. The NN’s parameters remain fixed, ensuring calibration only adjusts the output probabilities.
  2. 2.
    Isotonic Regression: This is a non-parametric method that learns a non-decreasing, piecewise-constant function $f(\cdot)$ to map uncalibrated probabilities $\hat{p}_{i}$ to calibrated values $\hat{q}_{i}=f(\hat{p}_{i})$, where $f(\cdot)$ is optimized to minimize the squared loss:
    $\sum_{i=1}^{n}\left(f(\hat{p}_{i})-y_{i}\right)^{2},$ (31)
    where $y_{i}$ denotes the true label. We implement this using scikit-learn’s IsotonicRegression, which fits the function using the pool-adjacent-violators algorithm. This ensures monotonicity while minimizing the squared error on the validation set. Once fitted, the model adjusts new predictions by applying the learned isotonic function to improve alignment with observed outcomes.
V Results and Demonstration
This section presents numerical results that demonstrate how the output of a calibrated predictor relates to the system’s OP and helps identify an appropriate $q_{th}$ to meet reliability targets. We begin by an evaluation of the predictor’s calibration using post-processing methods and reliability diagrams with logarithmic binning. This is followed by a robust simulation-based analysis that demonstrate the theoretical results established in Theorems 1 and 2. Finally, we present a proof of concept for the resource allocation technique introduced in Section II.
Figure 3: Histogram-based reliability diagrams of our predictor utilizing OLF and BCE for a 4-resource system with a rate threshold $\gamma_{th}$ of 0.5. Specifically, plots are presented for $\mathtt{SNR}$ values of (a) 2 dB and (b) 8 dB for predictors trained with OLF and (c) 2 dB and (d) 8 dB for predictors trained with BCE with various calibration methods.
V-A Results
Fig. 3 presents the histogram-based reliability plots using logarithmic binning for our outage predictor with a 4-resource system where $\gamma_{th}=0.5$ and $q_{th}=0.5$. Fig. 3 (a) and (b) show logarithmic reliability plots for predictors trained with OLF at an $\mathtt{SNR}$ of 2 dB and 8 dB, respectively, while Fig. 3 (c) and (d) present the corresponding logarithmic reliability plots for predictors trained with BCE. Furthermore, Table II presents the NLL and BS values for predictors trained with OLF and BCE, comparing the performance of calibrated and uncalibrated predictors across different $\mathtt{SNR}$ levels.
TABLE II: NLL and Brier Score (BS) Values for different $q_{th}$ and $\mathtt{SNR}$ Levels: Uncalibrated vs. Calibrated
SNR OLF BCE
NLL BS NLL BS
Uncalibrated Platt Isotonic Uncalibrated Platt Isotonic Uncalibrated Platt Isotonic Uncalibrated Platt Isotonic
2 dB 0.33 0.12 0.08 0.09 0.05 0.02 0.09 0.09 0.09 0.03 0.03 0.03
8 dB 0.18 0.05 0.03 0.06 0.03 0.01 0.05 0.04 0.04 0.01 0.01 0.01
Figure 4: Comparison of $\!P^{\star}_{\infty}(\gamma_{th},q_{th})$ and $\mathbb{E}[Q^{\star}\!\mid Q^{\star}\leq q_{th}]$ across different $q_{th}$ values, in a 4-resource system using OLF with $\gamma_{th}\!=2.0$ and $\mathtt{SNR}=6\mathrm{dB}$.
Figure 5: Comparison of $P$ and $\mathbb{E}[Q^{\star}]$ across different numbers of resources $\mathcal{R}$, using OLF with $\gamma_{th}=1.0$ and $\mathtt{SNR}=6~{}\mathrm{dB}$.
From Fig. 3, it is evident that the application of both the calibration methods significantly enhances the calibration performance of the outage predictor. This improvement is not only observed visually but also supported by the lower NLL and BS values achieved post calibration. To better understand their impact, we assess the effectiveness of each calibration method on the OLF-trained outage predictor below.
  1. 1.
    Platt Scaling: This method improves calibration, as shown in Fig. 3 (a) for $q_{th}=0.5$ and $\mathtt{SNR}=2~{}\mathrm{dB}$, where the grey bars represent the uncalibrated results and the dark blue bars represent the Platt scaled results. The calibration improvement is also reflected in the reduced NLL and BS values after calibration (see Table II), with NLL decreasing from 0.33 (uncalibrated) to 0.12 (Platt scaled) and BS dropping from 0.09 (uncalibrated) to 0.05 (Platt scaled). In general, this method demonstrates its effectiveness primarily at higher confidence levels, particularly within the range of $10^{-2}$ to $1$.
  2. 2.
    Isotonic Regression: This method provides significant calibration improvements through non-parametric adjustments to predicted probabilities. For example, in Fig. 3 (b), for $q_{th}=0.5$ and $\mathtt{SNR}=8~{}\mathrm{dB}$, the dark red bars align closely with the perfect calibration line, illustrating its effectiveness. This is further validated by a notable reduction in NLL from 0.18 (uncalibrated) to 0.03 (isotonic scaled) and in BS from 0.06 (uncalibrated) to 0.01 (isotonic scaled). It can be observed that this method proves to be effective even at lower confidence levels.
Figure 6: $P^{\star}_{|\mathcal{R}|}(\gamma_{th},q_{th})$ vs $q_{th}$ for a 4-resource system using OLF with (a) $\gamma_{th}~{}=0.5$, at $\mathtt{SNR}$ of 2 dB and (b) $\gamma_{th}=1.0$ at $\mathtt{SNR}$ of 8 dB.
Figure 7: $P^{\star}_{|\mathcal{R}|}(\gamma_{th},q_{th})$ vs $q_{th}$ for a 4-resource system using BCE with (a) $\gamma_{th}~{}=0.5$, at $\mathtt{SNR}$ of 2 dB and (b) $\gamma_{th}=1.0$ at $\mathtt{SNR}$ of 8 dB.
From Fig. 3 (c), and (d), it can be seen that the predictors trained with BCE display better calibration in their uncalibrated state compared to those trained with OLF. This is because BCE-based training is independent of $q_{th}$, enabling it to model the overall channel distribution. Post calibration, the predictors trained with BCE showed minimal changes, as reflected by the NLL values and BS (see Table II).
Figs. 4 and 5 illustrate the key results [^3] These plots do not compare the outage performance of OLF with BCE as $q_{th}$ is not a hyperparameter for BCE-based training. discussed in Section III, demonstrating the behaviour established for a perfectly calibrated predictor in Theorem 1. For this analysis, we apply isotonic regression, as it consistently achieves better calibration performance, according to the NLL and BS metrics, compared to Platt scaling. In particular, Fig. 4 shows the OP for an infinite resource limit, $P^{\star}_{\infty}(\gamma_{th},q_{th})$ alongside the expected confidence of the predictor when its output is below the threshold $q_{th}$, i.e., $\mathbb{E}[Q^{\star}\!\mid\!Q^{\star}\leq q_{th}]$. The figure compares calibrated and uncalibrated predictors over a range of $q_{th}$ values, using a system with 4 resources, $\gamma_{th}=2.0$, and $\mathtt{SNR}=6~{}\mathrm{dB}$. It can be observed that these curves align closely for the calibrated predictor when isotonic regression is applied, consistent with the behavior described in Theorem 1, Eq. (17).
Furthermore, this result helps in selecting an appropriate $q_{th}$ to meet a target system OP. For instance, a service provider might have a service level agreement (SLA) requiring the OP to be no greater than $0.01$ (i.e., a user should experience an outage no more than once every 100 transmissions). To satisfy this requirement, the provider can consult the calibrated outage predictor [^4] The NN used here is lightweight, enabling fast training, in contrast to large-scale models with extensive parameter counts.
This example applies only to a calibrated predictor, for which (17) holds. In the uncalibrated case, the output confidence does not reflect the true OP, this equation no longer holds, and $q_{th}$ must be chosen via empirical search.
. Given $\gamma_{th}=2.0$, $\mathtt{SNR}=6~{}\mathrm{dB}$, and a sufficient number of resources, Fig. 4 shows that setting $q_{th}=0.2$ with a well-calibrated predictor satisfies this SLA. Additionally, we observe that the calibrated and uncalibrated OLF-trained predictors both achieve the same minimum OP but at different $q_{th}$ values. The shift in the $q_{th}$ value where the calibrated predictor achieves its minimum OP is a consequence of calibration, which adjusts the predicted confidence scores to better reflect true OPs thereby rescaling the threshold at which the minimum OP occurs.
Fig. 5 compares the single-resource OP, $P$, with the expected predictor output, $\mathbb{E}[Q^{\star}]$, evaluated at $\gamma_{th}=1.0$ and ${}\mathtt{SNR}~{}=6~{}\mathrm{dB}$ for different number of resources. It can be seen that the calibrated predictor produces $\mathbb{E}[Q^{c}]$ values that closely match the observed single-resource OP $P$, consistent with Theorem 1, eq. (16).
Figs. 6 and 7 presents the OP for a 4-resource system using OLF and BCE, respectively, across a range of $q_{th}$ values, for (a) $\gamma_{th}=0.5$ at $\mathtt{SNR}=2~{}\mathrm{dB}$ and (b) $\gamma_{th}=1.0$ at $\mathtt{SNR}~{}=~{}8~{}\mathrm{dB}$. Unlike Fig. 5 from [46], these plots use $q_{th}$ values on a logarithmic scale and also present the OP obtained using the calibrated predictors. It can be seen that OP obtained using OLF and BCE exhibit a ‘U’ shaped curve, indicating that both very small and very large $q_{th}$ values are not good. This is because very small $q_{th}$ values results in the user almost always being allocated the final resource, while very large $q_{th}$ values lead to allocation to the first resource. Notably, at $q_{th}=0.5$, the uncalibrated OLF-trained predictor outperforms the BCE-trained ones. As $q_{th}$ decreases, this performance gap narrows, and in some cases, BCE slightly outperforms OLF at lower $q_{th}$.
Similar to the observation made in Fig. 4, it can also be seen from Fig. 6 that the calibrated and uncalibrated OLF-trained predictors both achieve the same minimum OP but at different $q_{th}$ values. For instance, in a 4-resource system with $\gamma_{th}=0.5$ and $\mathtt{SNR}=2~{}\mathrm{dB}$ (see Fig. 6 (a)), the uncalibrated predictor trained with OLF achieves its minimum OP at $q_{th}=0.5$, which matches its training threshold, while the calibrated predictor reaches the same minimum OP at $q_{th}=10^{-3}$. This observation aligns with the result of Theorem 2, which establishes that post-processing calibration cannot reduce the minimum achievable OP. It is important to note that this observation is not reflected in Fig.7, since $q_{th}$ is not a hyperparameter in BCE-based training. As a result, Theorem 2 is not applicable here.
V-B Proof of Concept
To demonstrate the application of the resource allocation technique and the OLF discussed in Section II[71] showcases video transmission over fading wireless channels using four different outage predictors. The goal is to compare the effectiveness of these predictors in reducing the likelihood of video outages and ensuring smooth playback. In this setup, a resource refers to a wireless channel used for video transmission. When a channel’s signal quality degrades due to fading and can no longer support reliable video transmission, an outage is likely to occur. However, the system leverages the outage predictor to anticipate such events and proactively switch to a more reliable channel. Playback is interrupted only when a predictor incorrectly recommends a channel that fails to meet the transmission requirement, in which case the screen turns red to indicate the outage (see Fig 8 (a)).
Figure 8: Snapshots from the video demonstration comparing a dummy model (non-ML strategy; top left) with ML-based outage predictors trained using BCE (top right), MSE (bottom left), and the OLF (bottom right). (a) Red frames indicate instances where the predictor incorrectly recommended a channel to support the video, while smooth playback reflects more effective channel selection. (b) The average outage rate over the entire video duration corresponding to each predictor.
For this demonstration, we set $\gamma_{th}=2.2$, and $\mathtt{SNR}$ = 15 $\mathrm{dB}$. The video runs at 25 frames per second, with predictions made 80 ms ahead, which is equivalent to two frames. The predictor is trained to identify whether an outage will occur within the next two frames, using either the BCE, MSE, or OLF. A baseline, non-ML strategy, referred to as a dummy model, is also included. This model always selects the first channel regardless of outage conditions.
This demonstration reinforces the effectiveness of the OLF-trained predictor, which resulted in an average outage rate of only 0.1% over the entire video duration, significantly lower than that of the the dummy model (6.8%), and the predictors trained using BCE (3.2%) and MSE (3.1%), leading to smoother video playback. Please note that the demonstration here uses the uncalibrated predictor with $q_{th}$ set to 0.5. However, it is worth highlighting that the calibrated predictor, trained using the OLF, will also achieve at best the same OP performance through the selection of an appropriate $q_{th}$, in accordance with Theorem 2 presented in this work. This behavior is consistent with Fig. 6, which shows that both the calibrated and uncalibrated predictors trained with OLF achieve the same minimum system OP, but at different $q_{th}$ (classification threshold) values.
VI Conclusion
This paper studies the calibration performance of the outage predictor used in an ML-assisted resource allocation system. A preliminary proof of concept demonstrating the application of the resource allocation technique referred to here, aimed at a broader audience, can be found in [71]. In this paper, we established key theoretical properties of the OP under perfect calibration. Specifically, we demonstrated that, in the infinite-resource case, the OP corresponded to the expected predictor output conditioned on it being below the classification threshold, while in the single-resource case, it equaled the predictor’s overall expected confidence. We then presented OP conditions for a perfectly calibrated predictor. These findings help guide the service provider in selecting $q_{th}$ values that satisfy the reliability requirements specified in an SLA. For instance, if an SLA permits no more than 1 outage in 200 transmissions (i.e., a system OP of $5\times 10^{-3}$), system engineers can rely on a well-calibrated predictor to meet this requirement by setting $q_{th}=0.09$ (see Fig. 4).
Notably, we showed that post-processing calibration cannot improve the system’s minimum achievable OP. Although both calibrated and uncalibrated predictors can achieve the same minimum OP, only calibrated predictors produce confidence scores that accurately reflect the true outage likelihood, enabling appropriate selection of $q_{th}$ values. In contrast, for uncalibrated predictors, $q_{th}$ must be tuned empirically. Furthermore, we identified a monotonicity-based condition on the predictor’s accuracy-confidence function that ensures improvement in the system’s OP, while also identifying a critical point where it remains unchanged.
Appendix A Appendix
A-A Proof of Theorem 1
We consider the proof of this theorem in three parts.
Before proceeding, we recall that the accuracy-confidence function associated with the predictor is defined as:
$A(q)\triangleq\mathbb{P}\left[C<\gamma_{th}\mid Q^{\star}=q\right],$ (32)
where perfect calibration corresponds to
$A(Q^{c})=Q^{c}.$ (33)
  1. 1.
    For a single-resource system: Recall from (2), the OP of a single resource system is given by:
    $P\triangleq\mathbb{P}\left[C<\gamma_{th}\right].$
    Using the law of total expectation:
    $P$ $=\mathbb{E}_{Q^{\star}}\left[\mathbb{P}\left[C<\gamma_{th}\mid Q^{\star}\right]\right]$ (34)
    $=\mathbb{E}_{Q^{\star}}[A(Q^{\star})].$ (35)
    For perfect calibration, we obtain
    $P=\mathbb{E}[Q^{c}],$ (36)
    where (36) follows from (33).
  2. 2.
    For ${|\mathcal{R}|}\rightarrow\infty$: The OP in the infinite resource limit is given by:
    $P_{\infty}(\gamma_{th},q_{th})$ $=\mathbb{P}\left[C<\gamma_{th}\mid Q^{\star}\leq q_{th}\right]$ (37)
    $=\mathbb{E}_{Q^{\star}\leq q_{th}}\left[\mathbb{P}\left[C<\gamma_{th}\mid Q^{\star}\right]\right]$ (38)
    $=\mathbb{E}_{Q^{\star}\leq q_{th}}[A(Q^{\star})].$ (39)
    For perfect calibration, we obtain
    $P^{c}_{\infty}(\gamma_{th},q_{th})=\mathbb{E}[Q^{c}\mid Q^{c}\leq q_{th}].$ (40)
    Here, (38) is given by the expectation taken over the predictor’s output $Q^{\star}\leq q_{th}$. Equation (39) follows from (32) and (40) follows from (33).
  3. 3.
    For an $|\mathcal{R}|$ resource system: Substituting (40) and (36) in  [46, eq. (13)], we obtain (18). This completes the proof.
A-B Proof of Theorem 2
Recall from (7) we have $Q^{c}=\;r\circ Q,$
$\implies$ $P^{c}_{|\mathcal{R}|}(\gamma_{th},q_{th})=\mathbb{P}\left(C<\gamma_{th}\mid Q^{c}\leq q_{th}\right),$ (41)
$\implies$ $\min_{q_{th}}P^{c}_{|\mathcal{R}|}(\gamma_{th},q_{th})\geq\min_{q_{th}}P_{|\mathcal{R}|}(\gamma_{th},q_{th}),$ (42)
where (42) follows from the fact that, minimizing OP over $Q^{c}$ corresponds to minimizing over a subset of the thresholds available through $Q$. This completes the proof.
A-C Proof of Theorem 3
We first show that $\mathbb{E}[A(Q^{\star})\mid Q^{\star}\leq q_{th}]<P$ implies $\mathbb{E}[A(Q^{\star})\mid Q^{\star}\leq q_{th}]<\mathbb{E}[A(Q^{\star})\mid Q^{\star}>q_{th}]$:
$\noindent\hskip 1.00006pt\mathbb{E}[A(Q^{\star})\mid Q^{\star}\leq q_{th}]$ $<P$
$=\mathbb{E}[A(Q^{\star})\mid Q^{\star}\leq q_{th}]F(q_{th})$
$+\mathbb{E}[A(Q^{\star})\mid Q^{\star}>q_{th}](1-F(q_{th}))$ (43)
[…]
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}<\mathbb{E}[A(Q^{\star})\mid Q^{\star}>q_{th}](1-F(q_{th}))$ (44)
[…]
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}<\mathbb{E}[A(Q^{\star})\mid Q^{\star}>q_{th}](1-F(q_{th}))$ (45)
[…]
where (43) follows from the law of total expectation, and (44) follows from rearrangement.
Next, we show that $\mathbb{E}[A(Q^{\star})\!\mid Q^{\star}\!\leq q_{th}]\!<\mathbb{E}[A(Q^{\star})\mid Q^{\star}>\!q_{th}]$ implies $\mathbb{E}[A(Q^{\star})\mid Q^{\star}\leq q_{th}]<P$:
$P$ $=\mathbb{E}[A(Q^{\star})]$
$=\mathbb{E}[A(Q^{\star})\mid Q^{\star}\leq q_{th}]F(q_{th})$
$~{}~{}~{}+\mathbb{E}[A(Q^{\star})\mid Q^{\star}>q_{th}](1-F(q_{th}))$ (47)
$>\mathbb{E}[A(Q^{\star})\mid Q^{\star}\leq q_{th}]F(q_{th})$
$~{}~{}~{}+\mathbb{E}[A(Q^{\star})\mid Q^{\star}\leq q_{th}](1-F(q_{th}))$ (48)
$=\mathbb{E}[A(Q^{\star})\mid Q^{\star}\leq q_{th}].$ (49)
Finally, we consider the critical point, defined by
$\mathbb{E}[A(Q^{\star})\mid Q^{\star}\leq q_{th}]=\mathbb{E}[A(Q^{\star})\mid Q^{\star}>q_{th}].$ (50)
Combining this with the law of total expectation, we have
$\mathbb{E}[A(Q^{\star})\mid Q^{\star}\leq q_{th}]=\mathbb{E}[A(Q^{\star})].$ (51)
Substituting this into (14) gives
$P^{\star}_{|\mathcal{R}|}(\gamma_{th},q_{th})$ $=\mathbb{E}[A(Q^{\star})](1-F(q_{th}))^{|\mathcal{R}|-1}$
$\hskip 1.00006pt~{}~{}~{}+\mathbb{E}[A(Q^{\star})]\left(1-(1-F(q_{th}))^{|\mathcal{R}|-1}\right)$
$=\mathbb{E}[A(Q^{\star})],$ (52)
meaning that the predictor yields no gain in the outage performance. This completes the proof.